HUBBLE SPACE TELESCOPE DAILY REPORT #3080 PERIOD COVERED: 0000Z (UTC) 03/25/02 - 0000Z (UTC) 03/26/02 Daily Status Report as of 085/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS/1 9581 (A BEA Science Program: A High Angular Resolution Survey of Some of the Galaxy's Most Massive Stars) Fine Guidance Sensor #1 was used to make optimal use of otherwise unused telescope time while HST's pointing is restricted to the BEA cone during the first 12 days of SMOV3B. The launch date of Feb 29 placed two of the Galaxy's most active star forming regions, Trumplers 14 and 16 in the Carina Nebula, within the BEA for the entire 12 day period. FGS1r was used to conduct a high angular resolution survey of the most massive and luminous stars in these two magnificent clusters to establish the incidence of binarity at the high mass end of the HR diagram. This survey will close the detection gap between spectroscopic and speckle surveys in which duplicity can be concealed and will dramatically improve our knowledge of the IMF at the high mass end of the main sequence as well as the luminosity function {LF} of star forming regions which produce such high mass stars. Finally, this survey will identify targets for an approved SIM Key Science project to dynamically determine O-star masses. There was no anomalous activity. 1.2 Completed Eighteen Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3 was used 1}.to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. As described in yesterday's report, the re-acquisition for the fifth iteration of this proposal failed, affecting that iteration when the take data flag remained down. Then, as further described in 2.1 and HSTAR 8579, during the eleventh iteration, the acquisition defaulted to fine lock backup on one FGS only, possibly affecting the observations in that iteration. Otherwise, no other problems were encountered. 1.3 Completed Two Sets of WF/PC-2 8954 (SM3B Point Spread Function Verification) The WF/PC-2 was used to verify the Point Spread Function after SM3B. There were no reported problems. 1.4 Completed Eight Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. As described in yesterday's report, the re-acquisition for the third iteration of this proposal failed, affecting that iteration when the take data flag remained down. Otherwise, there were no other problems reported. 1.5 Completed WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) The WF/PC-2 was used to monitor the planned cooldown to -88C, followed by UV monitors, focus checks, as well as decons and associated monitors. No problems were noted. 1.6 Completed STIS/CCD 9163 (Kinematics Of Emission-Line Gas Disks In Radio-Quiet Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to measure central black hole {BH} masses from the rotation rate of the emission-line gas, and to determine the nature and structure of the gas disks. Three galaxies with no radio jets and with Halpha+[NII] emission have been identified, which have dust disks similar to those commonly seen in our sample of radio-loud active galaxies. All observations completed nominally. 1.7 Completed Two Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.8 Completed Two Sets of FGS/1 8999 (SM3B Jitter Test) Fine Guidance Sensor 1R was used in both Position Mode {finelock} and Transfer Mode to observe selected targets in order to gain insight into the vehicle jitter characteristics after the alterations made to HST during SM3B. As further described in 2.1 and HSTAR 8579, during the second iteration, the acquisition defaulted to fine lock backup on one FGS only, possibly affecting the observations in that iteration. No other problems were encountered. 1.9 Completed Two Sets of STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.10 Completed STIS/CCD 9068 (Accurate Determination of the BH Mass in Early-Type Disk Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to map the ionized gas velocity field of three early-type disk galaxies, for which we will be able to derive high precision BH mass measurements. The sample galaxies have been selected by means of ground-based high resolution spectroscopy among 37 observed objects, since it is recognized in their central regions the clear presence of a circumnuclear Keplerian disk of ionized disk suitable for dynamical modelling. The proposal completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Per HSTAR 8579, the acquisition at 084/174019Z defaulted to fine lock backup on FGS-2 only. The proposals defined in1.2 and 1.8 may have been affected. Scheduled Re-acquisitions: 8 Successful: 8 2.2 FHST Updates: Scheduled: 15 Successful: 13 The full maneuver updates at 085/0612Z and at 085/0615Z failed due to high FHST vehicle errors on both trackers 1 and 3. HSTAR 8580 was written. 2.3 Operations Notes: Using ROP SR-9A, the SSR-3 EDAC error counter was cleared three times. The engineering status buffer limit for SESBSLD was modified four times per ROP DF-18A. During this operational day, the following operations notes were performed: Adjust ACS Error Count limit NICMOS MEB Limit Changes Turn Off ACS HRC/SBC Corrector Status limit NICMOS MEB Limit Changes NSSC-1 Load in Count Mode The SSR-3 playback pointer was set at 084/1240Z as directed by ROP SR-9A. A TTR was generated when there was a required re-transmit at 084/1413Z during a 486 uplink. Using ROP IC-1, SSA transmitter #2 was turned on for the interval 084/2257Z to 084/2324Z. The ephemeris table was uplinked at 085/0004Z per ROP DF-7A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW